# Losing Visual Needles in Image Haystacks: Vision Language Models are Easily Distracted in Short and Long Contexts

### v2

@@ Abstract
Aditya Sharma  Michael Saxon  William Yang Wang University of California, Santa Barbara LoCoVQA.github.io
11footnotetext: Equal contribution.
@@ 1 Introduction
Long-context vision-language models (VLMs) which accept multiple image inputs are pushing the boundaries of multimodal reasoning applications, but evaluation methods have not kept up. As open-weight long-context language models (LMs) have been around for the last few years Dong et al. (2023), researchers have developed robust evaluations to compare their capabilities against those of proprietary LMs Reid et al. (2024); Jiang et al. (2024a). However, until very recently Jiang et al. (2024b), the small set of long-context VLMs supporting multi-image input have been inaccessible to the public OpenAI et al. (2023), so evaluation practices for long-context VLMs Hong et al. (2023); Li et al. (2023b); Dai et al. (2024) are in their infancy Zhang et al. (2024). Most reasoning-based VLM evals have analogues in text-domain LM tasks. For example, image captioning Li et al. (2022); Wang et al. (2022); Nguyen et al. (2022) can be likened to text summarization, and visual question answering Chen et al. (2022); Wang et al. (2023); Xu et al. (2023a) to text QA.
@@
(a) 1-Image Context Length
(b) 4-Image Context Length
(c) 9-Image Context Length
Figure 2: Example of Image ($X$) corresponding to question-answer pair ($Q$, $A$) under increasing visual context lengths in the composed setting. The green box is for illustration purposes; not included in model inputs.
We aim to bridge gaps in multi-image VLM evaluation by drawing analogies to established long-context LM tasks. Long-document QA Fan et al. (2019) is an easy-to-evaluate test that directly showcases model performance on a useful application, and hints at its potential in retrieval-augmented generation (RAG) more broadly. Central to these evaluations is the model’s ability to recognize which elements in a lengthy input are necessary for answering a query and which are not, a skill we term extractive reasoning. Do long-context VLMs also exhibit these extractive reasoning capabilities? While current VLM benchmarks align with long-document summarization tasks, such as whole-video question answering Huang et al. (2021) or summarization Dilawari and Khan (2019), no existing benchmarks effectively capture a VLM’s ability to filter out irrelevant images in a long context to reason or answer a query—essentially, perform extractive reasoning over images.
@@ 2 Related Work
Text-based long-context tasks such as long-document question-answering (QA) Chen et al. (2023), summarization Phang et al. (2022), and retrieval-augmented generation (RAG) Lewis et al. (2020); Xu et al. (2023b) offer analogues to long-context reasoning in VLMs. Many of these tasks (e.g., QA, RAG) require language extractive reasoning, as discussed above.
@@
Figure 3: Each subfigure represents a variable number of hidden MNIST digits (1, 4, 8) in a 9-image context length.
@@ 3 LoCoVQA  Generation Method
Although this method can be applied to any vision-language dataset, we will describe the three specific tasks for which we generated long-context visual distractor benchmarks.
@@ 4 Experiments
Gemini 1.0 Pro Vision 32K Gemini
GPT-4 Vision 128K GPT-4
Gemini 1.5 Flash 1M Gemini
@@ 7 Discussion
However, the same may not necessarily be true for the sequential image language modeling task. In the existing interleaved vision-language training corpora, images are likely to followed immediately by their relevant text. As a result, attending to much earlier images in the documents may not be crucial for achieving low LM loss. Additionally, no open VLM is trained on image generation tasks conditioned on text while updating the core transformer weights. Overlooking this objective may prevent VLMs from robustly modeling the relationship between images and text, which could be crucial for performing well on these tasks.
@@ 8 Conclusion
Although LoCoVQA is a generalized process applicable to any VLM benchmark, we only evaluated it on three tasks. While the strong logarithmic decay trends between visual context length and performance observed across all three are compelling, expansion of LoCoVQA to additional tasks would render the findings clearer.
While LoCoVQA samples distractor images from the same datasets for open-domain and multiple-choice VQA, and our process appears to accurately filter out collisions, it is likely that a small number of collisions still occur, as it is inherently difficult to ensure no failures in an automated generating process Saxon et al. (2024). This may lead to a natural ceiling on VLM performance at each visual context length.
@@
Thanks to Yujie Lu, Luke Yoffe, Tejas Srinivasan, Weixi Feng, and Deepak Nathani for discussion and comments. This work was supported in part by the National Science Foundation Graduate Research Fellowship under Grant No. 1650114, and CAREER Award under Grant No. 2048122.
AS wrote all code, designed and executed experiments, produced all figures, conducted analysis, and wrote the manuscript. MS designed experiments, conducted analysis and was the primary writer and editor of the manuscript.
@@ Appendix A Evaluated Models
  1. (i)
    Moondream2-1.6b Moondream (2024) based on Phi-1.5 Li et al. (2023c),
  2. (ii)
    LLaVA-1.5 Liu et al. (2023a) with Vicuna-7b Peng et al. (2023) as the LLM backbone,
  3. (iii)
    LLaVA-1.6 (LLaVA-Next) Liu et al. (2024), an improvement over LLaVA-1.5 with higher image resolution and better visual reasoning, uses Mistral-7b Jiang et al. (2023),
  4. (iv)
    PaliGemma-3b Google (2024), based on open components from the SigLip Zhai et al. (2023) image encoder and the Gemma Team et al. (2024) language model,
  5. (v)
    Mantis-bakllava-7b Jiang et al. (2024b) fine-tuned from BakLLaVA SkunkworksAI (2023) and derived from LLaVA but using Mistral-7b Jiang et al. (2023),
  6. (vi)
    Mantis-Idefics2-8b Jiang et al. (2024b), the current state-of-the-art Mantis variant, based on Idefics2-8b Laurençon et al. (2024).
@@ Appendix D Model-level Scoring Details
A non-template response occurs when the model fails to enclose its output with <answer> tags. This is not penalized in our evaluations–where we still calculate exact match, BERT, ROUGE scores–it is crucial to monitor the trend. As visual context length increases VLMs often struggle to adhere to the prompted response formatting. Among the models tested, Gemini 1.5 demonstrates the best adherence to requested formatting, with only a single error. Gemini 1.0 exhibits slightly more issues, with no more than 100 errors in template adherence. GPT-4V shows the most significant struggle, with up to 600 formatting errors. We classify model refusals, possibly due to low confidence, as failures. Fundamentally, correct response formatting is necesary to enable use of LMs within
@@ Appendix E Supplementary Results
Figure 10 shows the radar plots for image context lengths $k=4,16,36$ to extend Figure 5. Table 2 and Table 3 display the $r^{2}$ values curve fits in Figure 4 fit. $p$-val asterisks denote the statistical significance of the overall fit. The red highlight signifies subchance performance (highlighted if more than half of the data points were below random choice). A down arrow indicates a negative correlation (as visual context length increase, performance increases).

### v3

@@ Abstract
[…]
@@ 1 Introduction
Long-context vision-language models (VLMs) which accept multiple image inputs are pushing the boundaries of multimodal reasoning applications, but evaluation methods have not kept up. As open-weight long-context language models (LMs) have been around for the last few years Dong et al. (2023), researchers have developed robust evaluations to compare their capabilities against those of proprietary LMs Reid et al. (2024); Jiang et al. (2024a). However, until very recently Jiang et al. (2024b), the small set of long-context VLMs supporting multi-image input have been inaccessible to the public OpenAI et al. (2023), so evaluation practices for long-context VLMs Hong et al. (2023); Li et al. (2023b); Dai et al. (2024) are in their infancy Zhang et al. (2024). Most reasoning-based VLM evals have analogues in text-domain LM tasks. For example, image captioning Li et al. (2022); Wang et al. (2022); Nguyen et al. (2022) can be likened to text summarization, and visual question answering (VQA) Chen et al. (2022); Wang et al. (2023); Xu et al. (2023a) to textual question answering (QA).
@@
(a) 1 image context
(b) 4 image composed context
(c) 9 image interleaved context
Figure 2: Example of Image ($X$) corresponding to question-answer pair ($Q$, $A$) under increasing visual context lengths in the composed and interleaved settings. The green box for illustration only, not included in model inputs.
We aim to bridge gaps in multi-image VLM evaluation by drawing analogies to established long-context LM tasks. Long-document QA Fan et al. (2019) is an easy-to-evaluate test that directly showcases model performance on a useful application, and hints at its potential in retrieval-augmented generation (RAG) more broadly. Central to these evaluations is the model’s ability to recognize which elements in a lengthy input are necessary for answering a query and which are not, a skill we dub extractive reasoning. Do long-context VLMs also exhibit these extractive reasoning capabilities?
While current VLM benchmarks align with long-document summarization tasks, such as whole-video question answering Huang et al. (2021) or summarization Dilawari and Khan (2019), no existing benchmarks effectively capture a VLM’s ability to filter out irrelevant images in a long context to reason or answer a query—essentially, perform extractive reasoning over images.
@@ 2 Related Work
Text-based long-context tasks such as long-document question-answering (QA) Chen et al. (2023), summarization Phang et al. (2022), and retrieval-augmented generation (RAG) Lewis et al. (2020); Xu et al. (2023b) offer analogues to long-context reasoning in VLMs. Many of these tasks (e.g., QA, RAG) require natural language extractive reasoning, as discussed above.
@@
Figure 3: Each subfigure represents a variable number of hidden MNIST digits in a 9 image composed context.
@@ 3 LoCoVQA  Generation Method
As a dynamic generator LoCoVQA  can convert any vision-language dataset into a long-context benchmark Saxon et al. (2024a). We describe the three datasets we used below.
@@ 4 Experiments
Gemini 1.0 Pro Vision 32K Gemini -
GPT-4 Vision 128K GPT-4 -
Gemini 1.5 Flash 1M Gemini -
@@ 7 Discussion
However, the same may not necessarily be true for the sequential image language modeling task. In the existing interleaved vision-language training corpora, images are likely to followed immediately by their relevant text. As a result, attending to much earlier images in the documents may not be crucial for achieving low LM loss. Additionally, none of the tested open VLMs are trained on image generation tasks conditioned on text while updating the core transformer weights. Overlooking this objective may prevent VLMs from robustly modeling the relationship between images and text, which could be crucial for performing well on these tasks.
@@ 8 Conclusion
Although LoCoVQA is a generalized process applicable to any VLM benchmark, we only evaluated it on three tasks. While the strong logarithmic decay trends between visual context length and performance observed across all three are compelling, expansion of LoCoVQA to additional tasks would render the findings clearer. While LoCoVQA samples distractor images from the same datasets for open-domain and multiple-choice VQA, and our process appears to accurately filter out collisions, a small number of collisions still occur—it is inherently difficult to ensure no failures in an automated generating process Saxon et al. (2024b). This may lead to a natural ceiling on VLM performance at each context length.
@@
Thanks to Yujie Lu, Luke Yoffe, Tejas Srinivasan, Weixi Feng, and Deepak Nathani for discussion and comments. This work was supported in part by the National Science Foundation Graduate Research Fellowship under Grant No. 1650114, and CAREER Award under Grant No. 2048122. AS wrote code, ran experiments, and produced figures. MS was the primary writer and editor of the manuscript. Both designed experiments and conducted analysis.
@@ Appendix A Evaluated Models
We evaluate: Moondream2-1.6b Moondream (2024) based on Phi-1.5 Li et al. (2023c), LLaVA-1.5 Liu et al. (2023a) with Vicuna-7b Peng et al. (2023) as the LLM backbone, LLaVA-1.6 (LLaVA-Next) Liu et al. (2024), an improvement over LLaVA-1.5 with higher image resolution and better visual reasoning, uses Mistral-7b Jiang et al. (2023), PaliGemma-3b Google (2024), based on open components from the SigLip Zhai et al. (2023) image encoder and the Gemma Team et al. (2024) language model, Mantis-bakllava-7b Jiang et al. (2024b) fine-tuned from BakLLaVA SkunkworksAI (2023) and derived from LLaVA but using Mistral-7b Jiang et al. (2023), Mantis-Idefics2-8b Jiang et al. (2024b), the current state-of-the-art Mantis variant, based on Idefics2-8b Laurençon et al. (2024).
@@ Appendix D Model-level Scoring Details
A non-template response occurs when the model fails to enclose its output with <answer> tags. This is not penalized in our evaluations–where we still calculate exact match, BERT, ROUGE scores–it is crucial to monitor the trend. As visual context length increases VLMs often struggle to adhere to the prompted response formatting. Among the models tested, Gemini 1.5 demonstrates the best adherence to requested formatting, with only a single error. Gemini 1.0 exhibits slightly more issues, with no more than 100 errors in template adherence. GPT-4V shows the most significant struggle, with up to 600 formatting errors. We classify model refusals, possibly due to low confidence, as failures.
@@ Appendix E Supplementary Results
Figure 10 extends Figure 5 for context lengths $k=4,16,36$. Table 2 and Table 3 display the $r^{2}$ values for Figure 4 fits, with $p$-val asterisks denoting statistical significance. Red highlighting signifies subchance performance (more than half of scores below random choice). Down arrows indicate positive correlation (improved performance with increasing context counters overall trend).
@@
Models Image Context 1st Quartile 2nd Quartile 3rd Quartile 4th Quartile Mantis $k=4$ 10.17% 10.73% 10.88% 10.96% $k=9$ 8.98% 10.28% 8.79% 10.04% $k=25$ 8.16% 8.21% 8.37% 8.64% $k=36$ 7.17% 7.72% 9.22% 7.75% GPT-4V $k=4$ 27.68% 23.03% 20.27% 23.90% $k=9$ 24.65% 15.96% 14.92% 18.92% $k=25$ 14.57% 10.51% 8.87% 10.12% $k=36$ 15.92% 9.06% 7.17% 6.96% Gemini 1.5 $k=4$ 33.28% 31.78% 33.28% 34.07% $k=9$ 28.37% 28.72% 29.93% 28.60% $k=25$ 25.69% 22.82% 23.07% 24.36% $k=36$ 23.09% 21.91% 21.91% 21.81%
Table 4: Impact of ordering on non-haystack tests—to assess if a positional bias akin to that in the needle in a haystack test is present for non-trivial tasks, we separately-averaged the OK-VQA scores for three models by the quartile the content image appears in the sequence. This confirms the same positional biases observed for MNIST haystack tests hold in the more challenging VQA setting.
Appendix F Case Study
List of Case Study Figures
  1. I
    OK-VQAExample 1
    Visual Context Length ($k=1$,   $k=4$,   $k=9$,   $k=25$,   $k=36$)
  2. II
    MMStar (SeedBench)Example 2
    Visual Context Length ($k=1$,   $k=4$,   $k=9$,   $k=25$,   $k=36$)
  3. III
    MMStar (MMBench)Example 3
    Visual Context Length ($k=1$,   $k=4$,   $k=9$,   $k=25$,   $k=36$)
  4. IV
    MMStar (MathVista)Example 4
    Visual Context Length ($k=1$,   $k=4$,   $k=9$,   $k=25$,   $k=36$)
Example 1 – OK-VQA
Question: What activity is this?
<image>
Visual Context Length ($k=1$)
   LLaVA 1.5 (cmp): This is a horseback riding activity
   LLaVA 1.6 (cmp): Horseback riding
   Moondream2 (cmp): Horseback riding
   Mantis (int): Horseback riding
   GPT-4V (cmp): horseback riding
   GPT-4V (int): horseback riding
   Gemini 1.0 (cmp): horseback riding
   Gemini 1.0 (int): horseback riding
   Gemini 1.5 (cmp): Horseback riding
   Gemini 1.5 (int): Horseback riding
   PaliGemma (cmp): horseback ride
Visual Context Length ($k=4$)
   LLaVA 1.5 (cmp): a girl riding a horse
   LLaVA 1.6 (cmp): Photography
   Moondream2 (cmp): biking
   Mantis (int): walking
   GPT-4V (cmp): horseback riding
   GPT-4V (int): horseback riding
   Gemini 1.0 (cmp): Horseback Riding
   Gemini 1.0 (int): Horseback Riding
   Gemini 1.5 (cmp): tennis
   Gemini 1.5 (int): tennis
   PaliGemma (cmp): horseback ride
Visual Context Length ($k=9$)
   LLaVA 1.5 (cmp): woman walking her horse
   LLaVA 1.6 (cmp): collage
   Moondream2 (cmp): biking
   Mantis (int): woman preparing to go surfing
   GPT-4V (cmp): bowling
   GPT-4V (int): horseback riding
   Gemini 1.0 (cmp): horseback riding
   Gemini 1.0 (int): horseback riding
   Gemini 1.5 (cmp): Bowling
   Gemini 1.5 (int): bowling
   PaliGemma (cmp): horseback riding
Visual Context Length ($k=16$)
   LLaVA 1.5 (cmp): This is a collage of various activities
   LLaVA 1.6 (cmp): Collage
   Moondream2 (cmp): Collage
   Mantis (int): a person riding a horse
   GPT-4V (cmp): surfing
   GPT-4V (int): Riding a horse
   Gemini 1.0 (cmp): surfing
   Gemini 1.0 (int): surfing
   Gemini 1.5 (cmp): horseback riding
   Gemini 1.5 (int): Surfing
   PaliGemma (cmp): collage
Visual Context Length ($k=25$)
   LLaVA 1.5 (cmp): This is a collage of photos, showcasing various activities.
   LLaVA 1.6 (cmp): Collage
   Moondream2 (cmp): Collage
   Mantis (int): Horseback riding
   GPT-4V (cmp): skiiing
   GPT-4V (int): playing catch with a frisbee
   Gemini 1.0 (cmp): This is a photo collage
   Gemini 1.0 (int): This is a photo collage
   Gemini 1.5 (cmp): horseback riding
   Gemini 1.5 (int): frisbee
   PaliGemma (cmp): skiing
Visual Context Length ($k=36$)
   LLaVA 1.5 (cmp): a man eating a piece of cake
   LLaVA 1.6 (cmp): motorcycle
   Moondream2 (cmp): colleage
   Mantis (int): small garden with a shed
   GPT-4V (cmp): surfing
   GPT-4V (int): Unable to provide the activities
   Gemini 1.0 (cmp): This is a game of Where’s Waldo.
   Gemini 1.0 (int): This is a game of Where’s Waldo.
   Gemini 1.5 (cmp): Surfing
   Gemini 1.5 (int): life
   PaliGemma (cmp): collage
Example 2 – MMStar (SeedBench)
Question: What is the color of the socks detected in the attribute detection of the <image>?
(A) Black (B) Red (C) White (D) Striped
<image>
Visual Context Length ($k=1$)
   LLaVA 1.5 (cmp): C
   LLaVA 1.6 (cmp): B
   Moondream2 (cmp): A
   Mantis (int): B
   GPT-4V (cmp): C
   GPT-4V (int): C
   Gemini 1.0 (cmp): B
   Gemini 1.0 (int): B
   Gemini 1.5 (cmp): A
   Gemini 1.5 (int): A
   PaliGemma (cmp): C
Visual Context Length ($k=4$)
   LLaVA 1.5 (cmp): A
   LLaVA 1.6 (cmp): B
   Moondream2 (cmp): A
   Mantis (int): B
   GPT-4V (cmp): C
   GPT-4V (int): C
   Gemini 1.0 (cmp): D
   Gemini 1.0 (int): C
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): C
   PaliGemma (cmp): A
Visual Context Length ($k=9$)
   LLaVA 1.5 (cmp): A
   LLaVA 1.6 (cmp): B
   Moondream2 (cmp): A
   Mantis (int): A: Black, B: Red, C: White
   GPT-4V (cmp): There are no visible socks in the image.
   GPT-4V (int): None of the images provided.
   Gemini 1.0 (cmp): C
   Gemini 1.0 (int): C
   Gemini 1.5 (cmp): Blue
   Gemini 1.5 (int): There are no socks in the image.
   PaliGemma (cmp): A
Visual Context Length ($k=16$)
   LLaVA 1.5 (cmp) A
   LLaVA 1.6 (cmp) B
   Moondream2 (cmp) A
   Mantis (int) A: Black, B: Red, C: White, D: Striped
   GPT-4V (cmp) There are no socks visible in the image
   GPT-4V (int) None of the above.
   Gemini 1.0 (cmp) C
   Gemini 1.0 (int) C
   Gemini 1.5 (cmp) C
   Gemini 1.5 (int) There are no socks in the image.
   PaliGemma (cmp) C
Visual Context Length ($k=25$)
   LLaVA 1.5 (cmp) C
   LLaVA 1.6 (cmp) B
   Moondream2 (cmp) A
   Mantis (int) B
   GPT-4V (cmp) C
   GPT-4V (int) Option C is not provided
   Gemini 1.0 (cmp) C
   Gemini 1.0 (int) C
   Gemini 1.5 (cmp) There are no socks in the image.
   Gemini 1.5 (int) There are no socks in the image.
   PaliGemma (cmp) B
Visual Context Length ($k=36$)
   LLaVA 1.5 (cmp) C
   LLaVA 1.6 (cmp) B
   Moondream2 (cmp) B
   Mantis (int) C
   GPT-4V (cmp) I am unable to provide assistance with this request
   GPT-4V (int) None of these
   Gemini 1.0 (cmp) A
   Gemini 1.0 (int) C
   Gemini 1.5 (cmp) There are no socks in the image.
   Gemini 1.5 (int) There are no socks in the image.
   PaliGemma (cmp) B
Example 3 – MMStar (MMBench)
Question: Which is the main topic of the <image>?
(A) A woman surfing (B) A man skating (C) A man surfing (D) A woman skating
<image>
Visual Context Length ($k=1$)
   LLaVA 1.5 (cmp): C
   LLaVA 1.6 (cmp): C
   Moondream2 (cmp): A
   Mantis (int): C
   GPT-4V (cmp): C
   GPT-4V (int): C
   Gemini 1.0 (cmp): C
   Gemini 1.0 (int): C
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): C
   PaliGemma (cmp): C
Visual Context Length ($k=4$)
   LLaVA 1.5 (cmp): C
   LLaVA 1.6 (cmp): C
   Moondream2 (cmp): A
   Mantis (int): C
   GPT-4V (cmp): C
   GPT-4V (int): C
   Gemini 1.0 (cmp): C
   Gemini 1.0 (int): C
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): C
   PaliGemma (cmp): C
Visual Context Length ($k=9$)
   LLaVA 1.5 (cmp): C
   LLaVA 1.6 (cmp): C
   Moondream2 (cmp): A
   Mantis (int): C: A man surfing, D: A woman skiting
   GPT-4V (cmp): C
   GPT-4V (int): -
   Gemini 1.0 (cmp): A
   Gemini 1.0 (int): C
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): C
   PaliGemma (cmp): C
Visual Context Length ($k=16$)
   LLaVA 1.5 (cmp): C
   LLaVA 1.6 (cmp): C
   Moondream2 (cmp): A
   Mantis (int): C: A man surfing, D: A woman skiting
   GPT-4V (cmp): C
   GPT-4V (int): -
   Gemini 1.0 (cmp): A
   Gemini 1.0 (int): A
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): C
   PaliGemma (cmp): C
Visual Context Length ($k=25$)
   LLaVA 1.5 (cmp): C
   LLaVA 1.6 (cmp): A
   Moondream2 (cmp): A
   Mantis (int): B
   GPT-4V (cmp): -
   GPT-4V (int): -
   Gemini 1.0 (cmp): A
   Gemini 1.0 (int): A
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): B
   PaliGemma (cmp): D
Visual Context Length ($k=36$)
   LLaVA 1.5 (cmp): C
   LLaVA 1.6 (cmp): A
   Moondream2 (cmp): A
   Mantis (int): D
   GPT-4V (cmp): C
   GPT-4V (int): -
   Gemini 1.0 (cmp): A
   Gemini 1.0 (int): A
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): C
   PaliGemma (cmp): C
Example 4 – MMStar (MathVista)
Question: What is the age gap between these two people in <image>? (Unit: years)
(A) 4 (B) 3 (C) 2 (D) 1
<image>
Visual Context Length ($k=1$)
   LLaVA 1.5 (cmp): E
   LLaVA 1.6 (cmp): A
   Moondream2 (cmp): A
   Mantis (int): A
   GPT-4V (cmp): I’m sorry, but I cannot provide information
   GPT-4V (int): I’m sorry, but I cannot provide information
   Gemini 1.0 (cmp): A
   Gemini 1.0 (int): A
   Gemini 1.5 (cmp): A
   Gemini 1.5 (int): A
   PaliGemma (cmp): A
Visual Context Length ($k=4$)
   LLaVA 1.5 (cmp): D
   LLaVA 1.6 (cmp): A
   Moondream2 (cmp): B
   Mantis (int): B
   GPT-4V (cmp): I’m sorry, but I cannot provide information
   GPT-4V (int): I’m sorry, but I cannot provide information
   Gemini 1.0 (cmp): B
   Gemini 1.0 (int): B
   Gemini 1.5 (cmp): D
   Gemini 1.5 (int): C
   PaliGemma (cmp): B
Visual Context Length ($k=9$)
   LLaVA 1.5 (cmp): E
   LLaVA 1.6 (cmp): A
   Moondream2 (cmp): B
   Mantis (int): B
   GPT-4V (cmp): I’m sorry, but I cannot provide information
   GPT-4V (int): I’m sorry, but I cannot provide information
   Gemini 1.0 (cmp): A
   Gemini 1.0 (int): D
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): C
   PaliGemma (cmp): B
Visual Context Length ($k=16$)
   LLaVA 1.5 (cmp): C
   LLaVA 1.6 (cmp): A
   Moondream2 (cmp): B
   Mantis (int): A
   GPT-4V (cmp): I’m sorry, but I cannot provide information
   GPT-4V (int): I’m sorry, but I cannot provide information
   Gemini 1.0 (cmp): B
   Gemini 1.0 (int): B
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): D
   PaliGemma (cmp): B
Visual Context Length ($k=25$)
   LLaVA 1.5 (cmp): E
   LLaVA 1.6 (cmp): A
   Moondream2 (cmp): B
   Mantis (int): B
   GPT-4V (cmp): I’m sorry, but I cannot provide information
   GPT-4V (int): -
   Gemini 1.0 (cmp): D
   Gemini 1.0 (int): D
   Gemini 1.5 (cmp): C
   Gemini 1.5 (int): C
   PaliGemma (cmp): A
Visual Context Length ($k=36$)
   LLaVA 1.5 (cmp): D
   LLaVA 1.6 (cmp): A
   Moondream2 (cmp): B
   Mantis (int): B
   GPT-4V (cmp): -
   GPT-4V (int): Impossible to determine
   Gemini 1.0 (cmp): A
   Gemini 1.0 (int): A
   Gemini 1.5 (cmp): D
   Gemini 1.5 (int): C
   PaliGemma (cmp): B